OpenStudy (anonymous):

solve for x 6x+2=x^2 +16x+18

jhonyy9 (jhonyy9):

\[x ^{2}+10x+16\]

OpenStudy (jamesj):

Right, x^2 + 10x + 16 = 0. Now, you should be able to see by inspection what are the factors. But if not, use the quadratic formula.

jhonyy9 (jhonyy9):

x_1,_2=(-16+/- sqrt(100-72))/2

OpenStudy (anonymous):

(x+8)(x+ 2)= 0

OpenStudy (anonymous):

Alboltz if you have trouble with factoring this may help x^2+10x+16=0 (x+a)(x+b)=0 you need to find a and b such that x^2+(a+b)x+(ab)=0 your coefficient in front of the x tern is 10 the coefficient in the last tern is 16 A+B=10 AB=16 now just find two numbers that satisfy both those equations (x+8)(x+2)=0 so your x's which satisfy this are x=-8 x=-2

OpenStudy (anonymous):

x^2+10x+16=0 (x+8)(x+2) x=-8, x=-2